OpenStudy (anonymous):

Photocopying. Quick Copy recently charged 49¢ per page for color copies and 7¢ per page for black-and-white copies. If Shirlee’s bill for 90 copies was $11.34, how many copies of each type were made?

OpenStudy (anonymous):

49x +7y =11.34 x+y= 90 solve the system :)

OpenStudy (anonymous):

i took 0.49 multipled is by 90 is that correct?

OpenStudy (anonymous):

because i got 44.1 +6.3=50.4 but im stumped after that

OpenStudy (anonymous):

yes sorry about that .49x +.07y = 11.34

OpenStudy (anonymous):

x = 90 - y substitute into the first equation .49 (90-y) + .07y = 11.34

OpenStudy (anonymous):

im so stumped because i got 50.4

OpenStudy (anonymous):

yes I am

OpenStudy (anonymous):

ty very much

OpenStudy (anonymous):

I apologize I had to leave earlier :( Here is the work. .49 (90-y) + .07y = 11.34 44.1-.49y +.07y=11.34 -.42y=-32.76 y=78 x=12